OpenStudy (anonymous):

Solve: 3^(2x+1) = 70

OpenStudy (anonymous):

is it \[3^{2x+1}=70 ?\]

OpenStudy (anonymous):

ya!

OpenStudy (anonymous):

take log on both sides 2x+1 log 3 = log 70

OpenStudy (anonymous):

2x+1 = log 70 / log 3

OpenStudy (anonymous):

you can solve it from here.